OpenStudy (anonymous):

Max is building a rectangular dog pen. One side of the pen will just be the side of the house. If he has 45 feet of fencing to use for the other three sides, what is the largest area he can create? Round to the nearest tenth.

OpenStudy (anonymous):

@imqwerty

OpenStudy (biohazard9064):

well what do you think you start with?

OpenStudy (anonymous):

The largest rectangular area is a square. Since the sides of a square are equal, just divide 45 by 3 to find the length of one side. Area of a square is side times side. Right ??

OpenStudy (biohazard9064):

yep

OpenStudy (anonymous):

But because it says round to the nearest 10th i thought it was wrong because it doesnt make sense .

OpenStudy (anonymous):

So its like 15x15 ??

OpenStudy (biohazard9064):

i believe so

OpenStudy (biohazard9064):

hold on

OpenStudy (anonymous):

Okay

OpenStudy (biohazard9064):

Yes I believe its 15 x 15 is it an answer choice or fill the blank?

OpenStudy (anonymous):

fill in the blank .

OpenStudy (biohazard9064):

So i would assume its 15 x 15 it makes sense but im gonna have @imqwerty double check

OpenStudy (anonymous):

Thank you

OpenStudy (mathmale):

Is this problem from algebra or from calculus? Your best bet would be to write the equation for the area of the rectangle. Because one side of the rectangle does not actually require fencing, it's not appropriate to assume that a square field would have the max area possible. Draw the field. Draw the 45 degree angle. Come up with a formula for the area of this field as a function of x. Graph this equation. It will likely rise, then fall, as x increases. Estimate the x value at which the area is at its max for this situation.

OpenStudy (anonymous):

Algebra

OpenStudy (mathmale):

Then drawing a graph of the area function would be your best bet.

imqwerty (imqwerty):

have been told this in class-> The largest rectangular area is a square. ? or you just took it yourself?

OpenStudy (anonymous):

I googled it

imqwerty (imqwerty):

okay we do it step by step :) you answer the question i ask

OpenStudy (anonymous):

Okay Cool

imqwerty (imqwerty):

okay we take it to be a rectangle of side lengths \(x\) and \(y\) now tell what will be the perimeter of this rectangle

OpenStudy (anonymous):

But i dont knoow the lengths ??

imqwerty (imqwerty):

yes we don't know but we assumed the side length to be \(x\) and \(y\) just write the perimeter in terms of \(x\) and \(y\)

OpenStudy (anonymous):

x+x +y+y ???

imqwerty (imqwerty):

yes correct :) so we can write it like this-> \(P=2x+2y\) okay?

OpenStudy (anonymous):

Okaay

imqwerty (imqwerty):

now try to isolate y :)

OpenStudy (anonymous):

do i just divide by 2 ?? or do i have to like move it to the other side first

OpenStudy (anonymous):

sorry im kinda stupid

OpenStudy (anonymous):

It's fifteen.

imqwerty (imqwerty):

no its not please refrain from giving direct answers :)

imqwerty (imqwerty):

okay to isolate y just do this-> \(2x+2y=P\) \(2x+2y-2x=P-2x\) \(2y=P-2x\) \(y=\large{\frac{P}{2}} -x\)

imqwerty (imqwerty):

okay? here we are trying to prove that rectangle of greatest area is a square with the help of an example..

OpenStudy (anonymous):

Yeah

imqwerty (imqwerty):

okay now tell what will be the Area of this rectangle

OpenStudy (mathmale):

The greatest possible area of a rectangle with four sides made up by fencing is A=x^2, where x is the length of one side. Yes. But if you need to form an area when one of the sides does not require fencing, the greatest possible enclosed area is NOT A=x^2.

OpenStudy (mathmale):

Again, I suggest you draw a picture and come up with a formula for the enclosed area. Then, graph this formula. At which x value do you obtain max area?

imqwerty (imqwerty):

yea drawing the graph is more easier approach :)

OpenStudy (mathmale):

Yes, AFTER you have developed a formula for the area in terms of one variable, x.

OpenStudy (anonymous):

Ill do that thanks guys .